Order: Lasix 0.1 mg/kg/hr as a continuous infusion

Available :10mg/mL

Clients weight : 68 kg

How many milliliters will the client receive every hour?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug client receive per hour.

       Volume of drug received

             Formula = Desired dose/Dose in hand * Quantity.

        Here

            Ordered dose = 0.1 mg/kg/hr.       

Weight of the patient = 68 kg.     

             Desired dose(D) = 0.1 * 68

                                         = 6.8 mg.

             Dose in hand (H) = 10 mg.

                     Quantity (Q) = 1 ml.

As per above formula

         Volume(ml) of drug required = D/H * Q

                                                      = 6.8/10 * 1

                                                      = 0.68 ml.

    Volume(ml) of drug receive every hour.

                Volume(ml) of drug = 0.68 ml.

                                Time (hr) = 1 hr.

                              ml/hr = 0.68 ml/hr.